OpenStudy (anonymous):

If 5 moles of P4 reacted with 22 moles Cl2 according to the above reaction, determine: a) How many moles PCl3 are produced a)________________ b) How many moles of P4 are left in excess after the reaction (if any) b)________________ c) How many moles of Cl2 are left in excess after the reaction (if any) c)________________

OpenStudy (anonymous):

Where is the reaction?

OpenStudy (anonymous):

You should first write and balance the equation. P4 + 6Cl2 - 4PCl3 Using ratio and proportion... According to our equation, 1 mol of P4 reacts with 6 of Cl2 and 1 mol of P4 as to 4 mol of PCl3 If you have 5 mol of P4, it should have reacted with 5*6/1=30 mol of Cl2 Or if you have 22 mol of Cl2, it should react with 22*1/6= 11/3 mol of P4 b) 5-11/3 = 1.33 a) 11/3 c) none, because Cl2 is the limiting reagent.